OpenStudy (anonymous):

@Loser66 @johnweldon1993 Cards numbered 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 are placed in a bag. If a card is drawn from the bag and then replaced 135 times, about how many times would a 6 be drawn? A. 2 B. 9 C. 15 D. 18

OpenStudy (anonymous):

Help please can some one explain how to do this

OpenStudy (johnweldon1993):

How many 6's are there? and how many TOTAL cards are there?

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

hi @beccaboo333 i mean Mrs.boo333

OpenStudy (anonymous):

so am i right 2

OpenStudy (johnweldon1993):

I was still waiting for you to answer the 2 questions I wrote above...

OpenStudy (anonymous):

@johnweldon1993 1.2 2.76

OpenStudy (anonymous):

am i right? @beccaboo333

OpenStudy (anonymous):

are u here @beccaboo333

OpenStudy (johnweldon1993):

76? O.o There are 2 6's in this data...and there are 15 total cards... so the probability of choosing a 6 would be \[\large \frac{2}{15}\] since each card is replaced 135 times we will have \[\large \frac{2}{15} \times 135 = \frac{270}{15} = ?\]

OpenStudy (anonymous):

u said total

OpenStudy (johnweldon1993):

Right....how many total cards are there.... 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 comes out to 15 cards...

OpenStudy (anonymous):

so C.

OpenStudy (anonymous):

Right?

OpenStudy (johnweldon1993):

No...The 15 was only used to figure out the probability of picking a '6' out of the total 15 cards... Look above to my previous post before that to see the explanation